Prove that the plane 2x-3y+6z=6 touches the conicoid 4x^2-9y^2+36z^2=36.
Find the point of contact.
1
Expert's answer
2019-03-25T12:03:10-0400

Solution:

Suppose that the plane


"2x-3y+6z=6"

touches the conicoid


"4x^2-9y^2+36z^2=36"

at some point


"M(x_0,y_0,z_0)."

Then the plane equation can be written as follows:


"{F_x}^,|_M(x-x_0)+{F_y}^,|_M(y-y_0)+{F_z}^,|_M(z-z_0)=0,"

where


"F(x,y,z)=4x^2-9y^2+36z^2-36,"


"{F_x}^,|_M=8x_0,{F_y}^,|_M=-18y_0,{F_z}^,|_M=72z_0;"

or


"8x_0(x-x_0)-18y_0(y-y_0)+72z_0(z-z_0)=0,"

"8x_0x-18y_0y+72z_0z-(8{x_0}^2-18{y_0}^2+{72z_0}^2)=0."

The coefficients of the last equation must be equal to the coefficients of the equation


"2x-3y+6z-6=0."

Check it out:


"8x_0=2;18y_0=3;72z_0=6;8{x_0}^2-18{y_0}^2+{72z_0}^2=6;"

or

"x_0=\frac{1}{4};y_0=\frac{1}{6};z_0=\frac{1}{12};\frac{1}{2}=6."

The last equality is wrong,so the plane


"2x-3y+6z=6"

does not touch the conicoid


"4x^2-9y^2+36z^2=36"

and there is no point of contact.

Answer:so the plane


"2x-3y+6z=6"


does not touch the conicoid


"4x^2-9y^2+36z^2=36"

and there is no point of contact.